OpenStudy (anonymous):

Sarah’s dad tells her she can have the greater of the following two amounts for her allowance (in dollars): The difference of 3 times 4 and 2 3 times the difference of 4 and 2 What is Sarah’s allowance, in dollars?

OpenStudy (texaschic101):

3(4) - 2 = 12 - 2 = 10 3(4 - 2) = 3(2) = 6 I think I did those right....you might want to get a second opinion.

OpenStudy (texaschic101):

that first equation is what I am having trouble with